solve the inequality
x^4-145x^2+144>_0

Answer by lwsshak3(11628)

solve the inequality
x^4-145x^2+144≥0
To solve this inequality, the left side must be factored first.
If you are allowed to use a graphics calculator, you will find the zeros at: -12, -1, 1, & 12.
If not, you will need to use the quadratic formula to find these zeros.
let u=x^2 and u^2=x^4
solve for u with following equation: u^2-145u+144=0
Then solve for x
..
(x+12)(x+1)(x-1)(x-12)≥0
number line:
<..+...-12...-...-1...+.....1....-..12....+......>
solution: (Interval notation)
(-∞,-12] U [-1,1] U [12,∞)